M= f^2 N divided by (n+N)^2

Solve for N. all variables represent all real numbers
Answer #1

SOLUTION :


M = f^2 N / (n + N)^2 

=> M(n + N)^2 = f^2 N

=> (n + N)^2 M = f^2 N


Comparing both sides :


(n + N)^2 = f^2. And N = M

=> n + N = f. And  N = M

=> N = f - n  and N = M. (ANSWER).

answered by: Tulsiram Garg